ira basketball team has 8 players, how many arrangements are there for the five different positions point guard, shooting guard, small forward, power forward, and center)?
Answer #1

Number of arrangements of 8 players taking 5 at a time =

P(8,5)=\frac{8!}{(8-5)!}=6720